K. Mello Company has three employees-a consultant, a computer programmer, and an administrator. The following payroll information is available for each employee: Consultant Computer Programmer Administrator Regular earnings rate $2,610 per week $32 per hour $46 per hour Overtime earnings rate Not applicable 2 times hourly rate 1.5 times hourly rate Federal income tax withheld $915 $255 $505 For hourly employees, overtime is paid for hours worked in excess of 40 hours per week. For the current pay period, the computer programmer worked 53 hours and the administrator worked 62 hours. Assume further that the social security tax rate was 6%, and the Medicare tax rate was 1.5%. Determine the gross pay and the net pay for each of the three employees for the current pay period. Assume the normal working hours in a week are 40 hours. If required, round your answers to two decimal places. Consultant Computer Programmer Administrator Gross pay $ 2,610 $ 82,752 $ Net pay $ $ $
Answer #1

Answer:

Consultant Gross pay = $2,610

Computer Programmer Gross Pay = 40 * 36 + (53 - 40) * 36 * 2 = $2,376

Administrator Gross pay = 40 * 46 + (62- 40) * 46 * 1.5 = $3,358